# ViTSGMM: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

### v1

@@ Abstract
We present ViTSGMM, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on STL-10 and CIFAR-10/100 datasets when using negligible labeled samples. Notably, this paper also reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results.  [^1] Code available at https://github.com/Shu1L0n9/ViTSGMM.
@@ 1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of semi-supervised learning is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing semi-supervised learning methods, such as consistency regularization and pseudo-labeling, rely on complex training techniques and architectural designs, and their generalization ability when the labeled datas are extremely limited remains to be improved. Furthermore, although some studies have attempted to apply the recently popular Vision Transformer (ViT) to semi-supervised learning, they have not yet fully leveraged its advantages.
@@
Gaussian Mixture Models (GMMs) are powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. In image recognition tasks, GMMs can model the distribution of data features and subsequently perform classification. The inspiration for this research comes from Miller & Uyar’s systematic study of the “Mixture of Experts” structure based on the joint probability model of features and labels. In our research, we found that the performance of GMMs is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
@@
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. Simultaneously, we incorporated semi-supervised learning strategies, utilizing a pseudo-labeling mechanism to efficiently leverage unlabeled data. This approach achieves superior classification performance within the framework of mixture density estimation. This method significantly overcomes the limitations of traditional methods in terms of feature representation capability and model generalization performance.
@@
Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised Gaussian mixture model (SGMM) method on three commonly used image classification datasets: STL-10, CIFAR-10, and CIFAR-100. The experimental results show that our method achieves state-of-the-art performance and exhibits good stability and robustness.
@@
Figure 1: Performance Comparison. Comparison of our method with recent semi-supervised learning models on STL-10 dataset with different numbers of labels, achieving state-of-the-art performance.
@@
It is worth mentioning that we have revealed a serious problem that has been overlooked by existing semi-supervised learning research: there are 7,500+ duplicate samples between the unlabeled samples of STL-10 and the test set samples, which is undoubtedly a serious data leakage problem when directly used for semi-supervised learning tasks.
@@
  • We introduce the powerful Vision Transformer (ViT) as a feature extractor into the classical Mixture of Experts (MoE) classifier for image recognition tasks.
  • We design a simple but performant semi-supervised learning strategy that effectively leverages negligible labels.
  • We demonstrate that this method achieves state-of-the-art performance.
  • We reveal the longstanding, yet overlooked issue of trainset and testset contamination in the STL-10 dataset, which has not been addressed in other works.
The remainder of this paper is organized as follows: Section 2 reviews related works; Section 3 introduces our ViTSGMM network; Section 4 demonstrates and analyzes experimental results; Section 5 discusses the significance and limitations of the research findings and outlines future research directions.
@@ 2 Related Works
2.1 Semi-Supervised Learning and Gaussian Mixture Model (GMM)
Semi-Supervised Learning (SSL) has been widely applied in tasks such as image classification and speech recognition in recent years, especially in scenarios where labeled data is scarce. Early SSL methods mainly relied on generating pseudo-labels and consistency regularization.
@@ 2.1 Semi-Supervised Learning and Gaussian Mixture Model (GMM)
Semi-Supervised Gaussian Mixture Model (SGMM) is the basic probabilistic framework of our method, combining GMM with unlabeled data, and proposes a learning method that simultaneously considers labeled data and unlabeled data by maximizing the joint log-likelihood function. Miller & Uyar (1996) proposed a classifier structure and learning algorithm that can effectively use unlabeled data to improve performance. There are also many impressive works such as Zong et al. (2018) using deep autoencoders and Gaussian mixture models for anomaly detection.
Most related to ours is the model of Zhao et al. (2023), which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to ViTSGMM, but it addresses the Generalized Category Discovery (GCD) problem from Vaze et al. (2022), aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with SGMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
2.2 Deep neural networks (DNN) and Visual Transformers (ViT)
Deep Neural Networks (DNNs) have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by Krizhevsky et al. (2012), in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of DNNs, many methods have been proposed to address overfitting issues, such as Dropout by Srivastava et al. (2014) and BatchNormalization by Ioffe & Szegedy (2015). The Residual Network (ResNet) proposed by He et al. (2016) effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by Yu et al. (2018) introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers (ViT) are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. Dosovitskiy et al. (2021) proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional CNN architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by Caron et al. (2021) learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
@@ 2.3 Self-Supervised Learning and Data Leakage Issues
In machine learning, the issue of data leakage has always been a potential challenge. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al. (2024) re-examines the task of data cleaning and formalizes it as a ranking problem.
@@ 3.2 Semi-supervised Mixture of Experts Classifier
Figure 2: Overview of our ViTSGMM network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled 1 represents a feature vector with a corresponding label, 2 represents a pseudo-labeled feature vector, and 3 represents an unlabeled feature vector.
@@ 3.3 Pseudo-labeling Mechanism
Algorithm 1 ViTSGMM Network
@@ 4.1 Standard SGMM Performance Analysis
Figure 3: Relationship between the number of labels and performance. The trend of accuracy with the number of labels for SGMM on the MNIST dataset shows a steady improvement in accuracy with an increase in labels.
@@
Figure 4: Relationship between PCA dimensions/Gaussian components and performance. The left figure shows the trend of SGMM accuracy with PCA dimensions on the MNIST dataset, while the right figure shows the trend of accuracy with the number of Gaussian submodels.
@@ 4.2 More Powerful Feature Extractors
Next, we performed t-SNE visualization on the features extracted by the three methods. In Figure 6, we present the t-SNE visualization results of feature representations obtained using different feature extraction methods on the CIFAR-10 dataset. It can be clearly seen from the figure that the feature points extracted by PCA exhibit a highly mixed state in the t-SNE space, and it is difficult to distinguish between samples from different categories. This indicates that the feature representation after PCA dimensionality reduction lacks sufficient discriminative power. In contrast, the feature points extracted using deep neural networks exhibit a certain clustering trend in the t-SNE space, but there is still some overlap between the boundaries of different categories. The feature points extracted using Visual Transformer (ViT) form a relatively clear cluster structure in the t-SNE space, which indicates that the feature representation extracted by ViT has stronger separability and can better capture the intrinsic structure of the data, providing a more favorable feature basis for subsequent classification tasks.
@@ 4.3 Experiment Setup
We report results on the Cifar-10/100 and STL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, with 10 classes in the training set, containing 5,000 labeled samples and 100,000 unlabeled images, where images have a resolution of 96x96 pixels. We performed special deduplication processing on the STL-10 dataset, remains 90,455 training images and 8,000 images for testing.
@@ 4.4 Comparison with State-of-the-art Methods
ViTSGMM(Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on STL-10.
In Table 2, we present the performance on STL-10. Our method achieves state-of-the-art performance on STL-10, outperforming other methods by a large margin. Our method achieves the best performance in all three settings, demonstrating the effectiveness of our proposed method. Compared to SequenceMatch, our ViTSGMM achieves +10.88%, +8.35%, and +1.41% on 40-label, 250-label, and 1000-label settings, respectively.
Table 2: Benchmark results on STL-10.
@@
In Table 3, we compare the performance of ViTSGMM with state-of-the-art methods on CIFAR-100 dataset. Our method outperforms other methods in the 40 labels setting, and achieves comparable performance in the 2500 labels and 10,000 labels settings.
@@
ViTSGMM(Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
@@ 4.5 Ablation Study
In order to test the effectiveness of the SGMM component, we compared it with the original classification head of ViT.
We trained on three datasets, error rates shown in Table 4, under scenarios with extremely few labels, the performance of SGMM is significantly superior to that of the Softmax layer.
@@
We conducted an ablation study to evaluate the impact of pseudo-labeling on the performance of ViTSGMM. We compared the performance of ViTSGMM with and without pseudo-labeling on three datasets, error rates shown in Table 4.
@@
Dataset CIFAR-10 CIFAR-100 STL-10
@@ 5 Conclusions
We introduce high-performance feature extractors into traditional generative models. Unlike traditional semi-supervised learning methods, this approach does not introduce domain-specific constraints or assumptions. The cross-modal unified modeling strategy performs surprisingly well when trained with a very small amount of labeled data and a large amount of unlabeled data. Extensive experiments show that ViTSGMM achieves new performance breakthroughs in benchmarks such as STL-10 and Cifar-10, maintaining over 96% classification accuracy even when using only 4 labeled samples per class.
Future development of this technology can proceed along several dimensions. First, constructing dynamic component adjustment mechanism based on a Bayesian framework to achieve parameter adaptivity. Secondly, enhancing the model’s discriminative performance when facing complex tasks such as fine-grained classification. Most interesting, exploring its application in other fields, such as electroencephalogram (EEG) signal analysis and medical image processing, thereby broadening its application scope and verifying its transferability. We believe these improvements will further unlock the technological benefits brought about by the fusion of probabilistic graphical models and deep learning.

### v2

# SemiOccam: A Robust Semi-Supervised Image Recognition Network Using Sparse Labels

@@ Abstract
We present SemiOccam, an image recognition network that leverages semi-supervised learning in a highly efficient manner. Existing works often rely on complex training techniques and architectures, requiring hundreds of GPU hours for training, while their generalization ability when dealing with extremely limited labeled data remains to be improved. To address these limitations, we construct a hierarchical mixture density classification decision mechanism by optimizing mutual information between feature representations and target classes, compressing redundant information while retaining crucial discriminative components. Experimental results demonstrate that our method achieves state-of-the-art performance on various datasets when using negligible labeled samples, and its simple architecture keeps training time to minute-level. Notably, this paper reveals a long-overlooked data leakage issue in the STL-10 dataset for semi-supervised learning tasks and removes duplicates to ensure the reliability of experimental results. We also release the deduplicated CleanSTL-10 dataset to facilitate fair and reliable research in future semi-supervised learning.  [^1] Code available at https://github.com/Shu1L0n9/SemiOccam.  [^2] CleanSTL-10 available at https://huggingface.co/datasets/Shu1L0n9/CleanSTL-10.
@@ 1 Introduction
Deep learning has achieved remarkable success in image classification tasks. However, its performance is highly dependent on large-scale annotated datasets. In real-world applications, labeled data is often scarce and expensive to acquire. Therefore, semi-supervised learning (SSL) has received widespread attention in recent years as an effective solution. The goal of semi-supervised learning is to train models using a small amount of labeled data and large amount of unlabeled data, thereby improving the model’s generalization ability. Nevertheless, most existing SSL methods, such as consistency regularization and pseudo-labeling, rely on complex training strategies and network designs, leading to training costs of hundreds of GPU hours even on high-performance computing resources like NVIDIA V100, and their generalization ability when labeled data are extremely limited still needs improvement. Moreover, although some recent works have explored Vision Transformers (ViT) in SSL, they have yet to fully exploit its advantages.
@@
Gaussian Mixture Models (GMMs) are powerful probabilistic models that assume data is generated from a mixture of multiple Gaussian distributions. The inspiration for this research comes from Miller & Uyar’s systematic study of the “Mixture of Experts (MoE)” structure based on the joint probability model of features and labels, as well as the principle of Occam’s razor.
According to Occam’s razor, among competing hypotheses that explain the data well, the simplest one is preferred. This motivates us to design a model that balances complexity and interpretability, ensuring efficient and effective feature representation without unnecessary overfitting. In our research, we found that the performance of GMMs is highly related to the quality of feature vectors. If the features lack representativeness or have insufficient discriminative power, the classification performance will be severely affected.
@@
To address this, we designed a feature encoding framework based on a self-attention mechanism to fully exploit the global contextual information in images and generate more discriminative feature representations. Simultaneously, we incorporated semi-supervised learning strategies, utilizing a pseudo-labeling mechanism to efficiently leverage unlabeled data.
@@
This approach achieves superior classification performance within the framework of mixture density estimation, significantly overcomes the limitations of traditional methods in terms of feature representation capability, training cost and model generalization performance. Taking the Dino with ViT-large backbone as our feature extraction model, we validated the proposed semi-supervised Gaussian mixture model (SGMM) method on various datasets, results show that our method achieves state-of-the-art performance within minutes on a single Tesla T4 GPU and exhibits good stability and robustness, with strong potential for edge deployment.
@@
Figure 1: Performance Comparison. Comparison of our method with recent semi-supervised learning models on datasets with different numbers of labels, achieving state-of-the-art performance with low training cost.
@@
It is worth mentioning that we have revealed a serious problem that has been overlooked by existing semi-supervised learning research: there are 7,500+ duplicate samples between the training samples and the testing samples in STL-10, which is undoubtedly a serious data leakage problem when directly used for semi-supervised learning tasks. To address this, we release the deduplicated CleanSTL-10 dataset for fair and reliable evaluation.
@@
  • We introduce the powerful ViT as a feature extractor into the classical yet often underappreciated MoE classifier for image recognition tasks.
  • We design a simple but performant semi-supervised learning strategy that effectively leverages negligible labels, achieves state-of-the-art performance.
  • We reveal the longstanding, yet overlooked issue of training set and test set contamination in the STL-10 dataset, which has not been addressed in other works.
  • We release the deduplicated CleanSTL-10 dataset to promote fair and reliable future research.
The remainder of this paper is organized as follows: Section 2 reviews related works; Section 3 introduces our SemiOccam network; Section 4 demonstrates and analyzes experimental results, including both classification performance and training cost; Section 5 discusses the significance and limitations of the research findings and outlines future research directions.
@@ 2 Related Works
2.1 Semi-Supervised Learning and Gaussian Mixture Models
Semi-Supervised Learning has been widely applied in tasks such as image classification and speech recognition in recent years, especially in scenarios where labeled data is scarce. Early SSL methods mainly relied on generating pseudo-labels and consistency regularization.
@@ 2.1 Semi-Supervised Learning and Gaussian Mixture Models
Semi-Supervised Gaussian Mixture Model is the basic probabilistic framework of our method. It combines GMM with unlabeled data and proposes a learning method that simultaneously considers both labeled and unlabeled data by maximizing the joint log-likelihood function. Our method is inspired by the classifier structure and learning algorithm proposed by Miller & Uyar, which effectively leverages unlabeled data to improve performance. In addition, many impressive works, such as Zong et al. (2018), have extended GMMs by incorporating deep autoencoders for tasks like anomaly detection.
Most related to ours is the model of Zhao et al. (2023), which also uses ViT for image embedding and then applies GMM for classification. This model is very similar to SemiOccam, but it addresses the Generalized Category Discovery (GCD) problem from Vaze et al. (2022), aiming to discover new categories, whereas our work goes further to demonstrate that combining deep learning with SGMM can outperform state-of-the-art methods, especially in scenarios with extremely limited labeled data, yielding significant advantages.
2.2 Deep neural networks and Visual Transformers
Deep Neural Networks have made significant progress in various supervised learning tasks in recent years. The breakthrough achieved by AlexNet, proposed by Krizhevsky et al. (2012), in image classification tasks laid the foundation for the development of deep learning. To improve the generalization ability of Deep Neural Networks (DNNs), many methods have been proposed to address overfitting issues, such as Dropout by Srivastava et al. (2014) and BatchNormalization by Ioffe & Szegedy (2015). The Residual Network (ResNet) proposed by He et al. (2016) effectively solved the gradient vanishing problem in deep network training by introducing residual connections, making it possible to train deeper networks. Subsequently, DLA by Yu et al. (2018) introduced a dynamic hierarchical attention mechanism on top of ResNet, further optimizing layer-wise interaction and feature fusion.
Visual Transformers are image classification models based on the Transformer architecture that capture global information in images through self-attention mechanisms. Dosovitskiy et al. (2021) proposed dividing images into fixed-size patches and applying self-attention mechanisms. In this way, ViT demonstrates strong performance in image classification tasks, especially on large-scale datasets, exhibiting higher generalization ability compared to traditional CNN architectures. The success of ViT has inspired many researchers to try applying it to self-supervised learning. The DINO method proposed by Caron et al. (2021) learns useful feature representations through self-supervised pre-training of ViT models. These representations can be fine-tuned on a small amount of labeled data to achieve high classification accuracy.
@@ 2.3 Self-Supervised Learning and Data Leakage Issues
Data leakage issue has always been a potential challenge in machine learning. Off-topic images, near-duplicate images, and labeling errors in benchmark datasets can lead to inaccurate estimations of model performance. Handling and removing duplicate samples has become an issue that cannot be ignored in semi-supervised learning. Therefore, Gröger et al. (2024) re-examines the task of data cleaning and formalizes it as a ranking problem.
@@ 3.2 Semi-supervised Mixture of Experts Classifier
Figure 2: Overview of our SemiOccam network. The network consists of a feature extractor and a semi-supervised mixture of experts classifier. An example of a sub-model is shown in the ellipse at the bottom right, where the circle labeled 1 represents a feature vector with a corresponding label, 2 represents a pseudo-labeled feature vector, and 3 represents an unlabeled feature vector.
@@ 3.3 Pseudo-labeling Mechanism
Algorithm 1 SemiOccam Network
@@ 4.1 Standard SGMM Performance Analysis
Figure 3: Relationship between number of labels/PCA dimensions/Gaussian components and performance on the MNIST dataset. The first figure shows a steady improvement in accuracy with an increase in labels. The second figure shows the trend of SGMM accuracy with PCA dimensions, while the right figure shows the trend of accuracy with the number of Gaussian submodels.
@@ 4.2 More Powerful Feature Extractors
Next, we performed t-SNE visualization on the features extracted by the three methods. In Figure 5, we present the t-SNE visualization results of feature representations obtained using different feature extraction methods on the CIFAR-10 dataset. It can be clearly seen from the figure that the feature points extracted by PCA exhibit a highly mixed state in the t-SNE space, and it is difficult to distinguish between samples from different categories. This indicates that the feature representation after PCA dimensionality reduction lacks sufficient discriminative power. In contrast, the feature points extracted using deep neural networks exhibit a certain clustering trend in the t-SNE space, but there is still some overlap between the boundaries of different categories. The feature points extracted using Visual Transformer form a relatively clear cluster structure in the t-SNE space, which indicates that the feature representation extracted by ViT has stronger separability and can better capture the intrinsic structure of the data, providing a more favorable feature basis for subsequent classification tasks.
@@ 4.3 Experiment Setup
We report results on the Cifar-10/100 and CleanSTL-10 datasets. CIFAR-10 contains 10 classes with 6,000 32x32 color images per class, totaling 60,000 images, of which 50,000 are for training and 10,000 for testing. CIFAR-100 is similar to CIFAR-10 but contains 100 classes with 600 images per class, of which 500 are for training and 100 for testing. STL-10 is an image classification dataset designed for self-supervised and semi-supervised learning, containing 10 classes with 5,000 labeled training samples and 100,000 unlabeled images, each at a resolution of 96×96 pixels. We created CleanSTL-10 by carefully removing duplicates and overlaps between training and test sets, as detailed in Appendix A. After this deduplication, CleanSTL-10 contains 90,455 training images and 8,000 testing images, ensuring more reliable and fair evaluation.
@@ 4.4 Comparison with State-of-the-art Methods
SemiOccam(Ours) 3.51 $\pm$ 0.12 3.47 $\pm$ 0.17 3.45 $\pm$ 0.16
Results on CleanSTL-10.
In Table 2, we present the performance on CleanSTL-10. Our method achieves state-of-the-art performance, outperforming other methods by a large margin. Our method achieves the best performance in all three settings, demonstrating the effectiveness of our proposed method. Compared to SequenceMatch in 2024a, our SemiOccam achieves +10.88%, +8.35%, and +1.41% on 40-label, 250-label, and 1000-label settings, respectively.
Table 2: Benchmark results on CleanSTL-10.
@@
In Table 3, we compare the performance of SemiOccam with state-of-the-art methods on CIFAR-100 dataset. Our method outperforms other methods in the 400 labels and 2,500 labels setting, and achieves comparable 10,000 labels settings.
@@
SemiOccam(Ours) 26.59 $\pm$ 1.02 22.19 $\pm$ 0.81 21.21 $\pm$ 0.26
4.5 Training Cost Analysis
In the field of semi-supervised learning, besides model performance, training overhead is a key metric for evaluating the practicality and environmental impact of algorithms. This section will quantitatively compare the training overhead of existing benchmarks with our proposed method.
TorchSSL from Zhang et al. is a mainstream SSL evaluation protocols, typically employ training deep neural networks from scratch, leading to significant computational burdens. For example, training the FixMatch algorithm on a single NVIDIA V100 GPU requires “300 hours $\times$ 3 settings $\times$ 3 seeds” for the CIFAR-100 dataset; “110 hours $\times$ 3 settings $\times$ 3 seeds” for CIFAR-10; and “225 hours $\times$ 3 settings $\times$ 3 seeds” for STL-10. The total training time amounts to 5715 GPU hours (approximately 238 GPU days). This high computational demand poses a significant experimental cost barrier for academia and resource-constrained research institutions.
USB from Wang et al. is another current mainstream SSL evaluation protocol. It reports that training the FixMatch algorithm on a single NVIDIA V100 GPU incurs an overhead of “11 hours $\times$ 2 settings $\times$ 3 seeds” for the CIFAR-100 dataset, and “18 hours $\times$ 2 settings $\times$ 3 seeds” for the STL-10 dataset. Across these two datasets, a total of 174 GPU hours are required.
Our proposed method utilizes a pre-trained ViT as a feature extractor. After this feature extraction step is completed, the core SGMM classifier training process can be completed in less than ten minutes on a free Tesla T4 GPU 15GB on the Google Colab platform, demonstrating high efficiency and state-of-the-art performance. This minute-level training time gives our method a significant competitive advantage in resource-constrained practical deployments and rapid iteration research scenarios.
@@ 4.6 Ablation Study
In order to test the effectiveness of the SGMM component, we compared it with the original classification head of ViT. We trained on three datasets, error rates shown in Table 4, under scenarios with extremely few labels, the performance of SGMM is significantly superior to that of the Softmax layer.
@@
We conducted an ablation study to evaluate the impact of pseudo-labeling on the performance of SemiOccam. We compared the performance of SemiOccam with and without pseudo-labeling on three datasets, error rates shown in Table 4.
@@
Dataset CIFAR-10 CIFAR-100 CleanSTL-10
@@ 5 Conclusions
We introduce high-performance feature extractors into the classical yet often underappreciated MoE classifier for image recognition tasks. Our architecture is simple and training is highly efficient, making it particularly well-suited for scenarios with limited labeled data. The proposed unified modeling strategy performs surprisingly well when trained with very few labeled samples and large-scale unlabeled data. In our study, we also identify a long-overlooked data leakage issue in the STL-10 dataset and release the deduplicated CleanSTL-10 dataset to ensure fair and reliable evaluation in semi-supervised learning. Extensive experiments show that SemiOccam achieves new performance breakthroughs on benchmarks such as CleanSTL-10 and CIFAR-10, maintaining over 96% classification accuracy even with only 4 labeled samples per class in under ten minutes on a single Tesla T4 GPU, compared to the hundreds of GPU hours required by existing methods.
Future development of this technology can proceed along several dimensions. First, constructing dynamic component adjustment mechanism based on a Bayesian framework to achieve parameter adaptivity. Secondly, enhancing the model’s discriminative performance when facing complex tasks such as fine-grained classification. Most interesting, exploring its application in other fields, such as electroencephalogram signal analysis and medical image processing, thereby broadening its application scope and verifying its transferability.
We believe these improvements will further unlock the technological benefits brought about by the fusion of probabilistic graphical models and deep learning. Crucially, given our method’s advantages in short training times and high efficiency, we are particularly keen to see its widespread adoption in edge deployment scenarios and applications requiring rapid iteration.